OpenStudy (anonymous):

Which circles have a circumference of 24π units? Choose all answers that are correct. A. a circle with a radius of 6 units B. a circle with a radius of 12 units C. a circle with a diameter of 12 units D. a circle with a diameter of 24 units

OpenStudy (k_lynn):

The circumference formula is \[2\Pi r\] But since there is pi in the answer, just multiply like this \[2r\] R is the radius. Plug in all the radii given and see if you get 24.

OpenStudy (anonymous):

what. i have no clue what you just said. o-o

geerky42 (geerky42):

\(2\pi r\)*

OpenStudy (anonymous):

nope nah i dont know. I cant do math

geerky42 (geerky42):

\[C=2\pi r\quad\Longrightarrow\quad\dfrac{C}{2\pi}=r\] So you just divide circumference by \(2\pi\). What do you get?

OpenStudy (anonymous):

I dont know . e.e i cant do this shlit

geerky42 (geerky42):

You can... Circumference is \(24\pi\), so you have \(\dfrac{24\pi}{2\pi}\) What to do from here?

geerky42 (geerky42):

Do numerator and denominator have something in common that we can cancel out?